Yuri has a board that is 98in. long. He wishes to cut the board into two pieces so that one piece will be 10in. longer than the other. What should the length of each piece be?
Answer by checkley71(8403)

X+(X+10)=98
X+X=10=98
2X=98-10
2X=88
X=44 FOR THE SHORT PIECE.
44+10=55 FOR THE LONGER PIECE.
PROOF
44+54=98
98=98